### v2

# A Hybrid Multi-Factor Network with Dynamic Sequence Modeling for Early Warning of Intraoperative Hypotension

Abstract
Intraoperative hypotension (IOH) prediction using past physiological signals is crucial, as IOH can lead to inadequate organ perfusion, increasing the risk of severe complications and mortality. However, existing IOH prediction methods often rely on static modeling, overlooking the complex temporal dependencies and non-stationary nature of physiological signals. In this paper, we propose a Hybrid Multi-Factor (HMF) network that models IOH prediction as a dynamic sequence forecasting problem, explicitly capturing temporal dependencies and physiological non-stationarity.. Specifically, we formalize physiological signal dynamics as a sequence of multivariate time series, and decompose them into trend and seasonal components, enabling distinct modeling of long-term and periodic variations. For each component, we employ a patch-based Transformer encoder to extract representative features with the concern of computational efficiency and representation quality. Furthermore, to mitigate distributional drift arising from the evolving signals, we introduce a symmetric normalization mechanism. Extensive experiments on both a publicly available dataset and a private dataset collected from real-world hospital settings demonstrate that our approach significantly outperforms competitive baselines. We hope HMF offers a new perspective on IOH prediction and further enhances surgical safety [^2] https://github.com/Mingyue-Cheng/HMF.
1 Introduction
Figure 1: Illustration of MAP Sequence: Temporal Dynamics and Distribution Shifts in Intraoperative Hypotension Event.
Intraoperative mortality has decreased by a factor of $100$ over the past century, making deaths during surgery a rare occurrence Li et al. (2009); Xue et al. (2022). However, mortality within the first month following surgery remains a significant concern, with approximately 2% of patients undergoing inpatient noncardiac surgery dying within $30$ days postoperatively Spence et al. (2019)—amounting to more than $4$ million deaths worldwide each year Saugel and Sessler (2021). These postoperative deaths are most strongly linked to complications, which are often triggered by intraoperative events. Among these, intraoperative hypotension (IOH) is a common and serious complication Kim et al. (2023), characterized by a significant drop in Mean Arterial Pressure (MAP) [^3] MAP, derived from Arterial Blood Pressure (ABP), reflects tissue perfusion and is calculated as $MAP=DBP+\frac{1}{3}(SBP-DBP)$, where SBP and DBP denote peak and minimum arterial pressures per cardiac cycle. A sustained decline in SBP and DBP reduces MAP, risking organ hypoperfusion. sustained over a period of time.
Previous studies Hatib et al. (2018a) have shown that IOH events can be predicted using machine learning applied to physiological signals. Clinical trials further indicate that early warnings and timely interventions help mitigate hypotension severity and reduce postoperative complications Hwang et al. (2023); Fernandes et al. (2021). Recent advances in IOH prediction primarily follow a feature extraction-based approach Lee et al. (2021); Davies et al. (2020), where handcrafted features are used with classifiers like logistic regression and random forests. However, these methods face two key limitations: (1) reliance on handcrafted features limits their ability to capture temporal dependencies, and (2) their models lack the flexibility to adapt to complex real-world IOH scenarios. Additionally, IOH definitions vary across clinical settings, and fixed threshold-based criteria may fail to accurately characterize IOH, particularly in hypertensive or elderly patients, where more nuanced definitions are often needed.
To overcome the limitations of existing approaches, we reformulate IOH prediction as a multivariate time series forecasting problem. The core rationale behind this formulation is that time series modeling effectively captures the temporal dynamics and distribution shifts inherent in physiological signals. Since hypotensive events are defined based on MAP series, accurate sequence forecasting directly impacts risk identification, offering greater flexibility. However, as illustrated in Figure 1, several challenges arise in MAP series prediction. First, the temporal dynamic waveform of MAP consists of multiple components, including trends and periodic patterns, which may need to be modeled separately to improve forecasting accuracy. Second, the high information redundancy caused by the extended length of MAP waveforms increases computational complexity and complicates representation learning in sequence modeling. Finally, evolving distribution shifts introduce additional challenges, as sudden fluctuations in blood pressure alter the statistical properties (e.g., mean and standard deviation) of the MAP series, making it difficult to maintain consistent predictive performance. Addressing these challenges is essential for developing a high-accurate and adaptable IOH prediction framework.
To address these limitations, we introduce the Hybrid Multi-Factor (HMF) framework, which explicitly models temporal dependencies and mitigates distribution shifts, ensuring adaptive and accurate IOH prediction. A key feature of HMF is its use of sequence decomposition techniques to model the trend and seasonal components separately, allowing for a more structured representation of historical dynamics. By integrating multi-factor physiological signals, including MAP and SBP, HMF effectively captures both local fluctuations and long-term dependencies. To efficiently extract informative features, we adopt a patch-based Transformer encoder to learn representations of the given sequence. Additionally, a symmetric normalization mechanism is introduced to mitigate distribution shifts caused by evolving physiological signals. We conduct extensive experiments on both public and private datasets collected from real-world clinical scenarios. The results demonstrate the superiority of HMF in improving predictive performance while effectively modeling multi-factor temporal interactions.
The key contributions are summarized as follows:
  • We reformulate IOH prediction as a multivariate time series forecasting problem and introduce HMF, which explicitly captures physiological signal dynamics.
  • We highlight the challenges of IOH prediction based on historical physiological signal series, particularly the complexities in series structure and the non-stationary nature of physiological signals.
  • We validate the effectiveness of the proposed approach through extensive experiments on two real-world datasets, demonstrating superior performance compared to existing methods.
Figure 2: Feature Extraction and Temporal Window Setup for Blood Pressure Trend Prediction: Context, Skipped, and Target Windows.
Figure 3: Proposed Hybrid Multi-Factor Model Architecture for Intraoperative Hypotension Prediction.
2 Preliminaries
To formulate the IOH prediction task, we first define how physiological features are extracted and structured, which serves as the foundation for subsequent problem formulation.
2.1 Multi-factor Feature Construction
To predict intraoperative hypotension (IOH), real-time monitoring and modeling techniques analyze arterial blood pressure (ABP) dynamics. Identifying early patterns preceding hypotensive events is crucial for timely intervention. This study structures multi-factor ABP-derived features to enhance forecasting accuracy.
During surgery, ABP is continuously monitored at high frequency (e.g., 100 Hz), capturing hemodynamic fluctuations. Prior studies reveal a strong correlation between ABP variations within a context window and subsequent hypotension. Clinically, hypotension is defined based on Mean Arterial Pressure (MAP), the average pressure per cardiac cycle, considered hypotensive when sustained below a threshold (e.g., one minute). Systolic Blood Pressure (SBP), the peak pressure per cycle, is included due to its physiological link to MAP. These ABP-derived signals—MAP and SBP—serve as core predictive features.
To construct informative representations, we extract continuous ABP segments, removing outliers and artifacts. A sliding window approach processes ABP series, extracting SBP (peak pressure) and Diastolic Blood Pressure (DBP, lowest pressure). MAP is computed as:
$\text{MAP}=\frac{1}{3}\times\text{SBP}+\frac{2}{3}\times\text{DBP}.$ (1)
Hypotensive events are identified when MAP remains below 65 mmHg for at least one minute, following clinical guidelines. Event annotation details, including threshold selection and temporal alignment, are further provided in the experimental and supplementary material.
For IOH prediction, we construct a multivariate time series dataset incorporating SBP and MAP to capture both short-term fluctuations and long-term trends. This formulation enables effective feature extraction and temporal alignment, as illustrated in Figure 2. By integrating multiple ABP-derived factors, our approach improves the model’s ability to learn temporal dependencies and physiological dynamics associated with IOH.
2.2 Problem Definition
Building on the multi-factor feature construction framework, we formulate IOH risk prediction as a dynamic sequence forecasting problem. Given real-time arterial blood pressure (ABP) monitoring, our goal is to predict future mean arterial pressure (MAP) trends, enabling early warnings for potential hypotensive events. Formally, let $\mathbf{X}_{\text{MAP}}=\{x_{\text{MAP}}^{1},x_{\text{MAP}}^{2},\dots,x_{\text{MAP}}^{T}\}$ and $\mathbf{X}_{\text{SBP}}=\{x_{\text{SBP}}^{1},x_{\text{SBP}}^{2},\dots,x_{\text{SBP}}^{T}\}$ denote the historical MAP and systolic blood pressure (SBP) sequences over a context window of length $T$. The objective is to forecast the future MAP values $\mathbf{\hat{Y}}_{\text{MAP}}=\{\hat{y}_{\text{MAP}}^{T+1},\dots,\hat{y}_{\text{MAP}}^{T+\tau}\}$ over a prediction horizon $\tau$. A hypotensive event occurs if the predicted MAP falls below a predefined threshold $\theta_{\text{MAP}}$. Further details on event annotation and threshold selection are provided in supplement material part. By leveraging multi-factor temporal dependencies, our formulation enhances the reliability of IOH risk assessment, facilitating timely clinical intervention.
3 The Proposed HMF
This section provides a detailed description of the proposed HMF framework.
3.1 Framework Overview
The proposed Hybrid Multi-Factor (HMF) model predicts intraoperative hypotension (IOH) by leveraging MAP and SBP time series to capture both long-term trends and short-term fluctuations. As shown in Figure 3, the model consists of four key components: symmetric normalization, sequence decomposition, dynamic dependence modeling, and prediction. First, symmetric normalization reduces inter-patient variability. Sequence decomposition separates trend and seasonal components to enhance feature learning. Transformer encoders process each component independently to capture temporal dependencies. Finally, the model fuses learned representations and reconstructs the MAP trajectory for forecasting. By integrating multi-factor signals and sequence decomposition, the HMF model improves IOH prediction accuracy and adaptability to dynamic MAP variations.
3.2 Symmetric Normalization
Non-stationarity in MAP and SBP series, driven by physiological variations during surgery, causes sudden shifts in statistical properties like mean and standard deviation Kim et al. (2021), challenging consistent IOH prediction. To mitigate these shifts and enhance reliability, we introduce a symmetric normalization module that ensures consistent feature scales across different surgical conditions, stabilizing model training and improving generalization in unseen cases.
Given a context window $X\in\mathbb{R}^{L\times 2}$, instance normalization is applied as:
$X^{\prime}=\frac{X-\mu_{X}}{\sigma_{X}},$ (2)
where $\mu_{X}$ and $\sigma_{X}$ are the mean and standard deviation of $X$. This transformation allows the model to focus on relative changes rather than absolute values, reducing the impact of abrupt fluctuations and improving pattern recognition. Additionally, by normalizing across instances, the method mitigates inter-patient variability, ensuring stable feature distributions.
After prediction, de-normalization restores the original scale:
$\hat{Y}=\hat{Y}^{\prime}\cdot\sigma_{X}+\mu_{X},$ (3)
where $\hat{Y}^{\prime}$ is the predicted sequence and $\hat{Y}$ is the final output. This ensures stability while preserving raw characteristics, facilitating reliable and interpretable IOH forecasting.
3.3 Sequence Decomposition Preprocessing
The Arterial Blood Pressure(ABP) signal exhibits a complex waveform structure, consisting of multiple components like trend and periodic elements, which need to be individually modeled for accurate forecasting. After applying instance normalization, which results in the normalized sequence $X^{\prime}$, it becomes essential to decouple the physiological series into its trend and seasonal components. This decomposition allows for more precise modeling of each component, improving the overall forecasting accuracy. We follow the methodology of previous workWu et al. (2021) to achieve this decoupling, enabling the model to better capture the inherent patterns in the ABP signal, is computed using:
$\text{Trend}=\text{AvgPool}(\text{Padding}(X^{\prime})),$ (4)
$\text{Seasonal}=X^{\prime}-\text{Trend},$ (5)
where AvgPool smooths the sequence by downsampling through averaging within a specified window, while Padding ensures full coverage across the entire sequence.
3.4 Dynamic Feature Representation Layer
Patch Embedding.
To model the intricate dynamics of MAP and SBP series during surgery, we apply a patch embedding technique after the decomposition layer. This reduces sequence length, lowering computational complexity and improving efficiency. Patch-based modeling captures local patterns while mitigating noise and outliers, enhancing prediction robustness.
The decomposed series $W\in\mathbb{R}^{L\times 2}$ is transformed into a compact representation $W_{\text{patch}}\in\mathbb{R}^{\frac{L}{S}\times d_{\text{model}}}$ using three 1D convolutional layers:
$W_{\text{patch}}=\text{Conv1D}_{3}(W),$ (6)
[…]
To encode temporal structures, we introduce learnable positional encodings:
$W_{\text{pos}}=W_{\text{patch}}+\text{PositionalEncoding}(W_{\text{patch}}).$ (7)
These encodings help preserve temporal dependencies, crucial for MAP forecasting and IOH detection, where event timing directly impacts prediction accuracy.
Sequence Dependence Modeling.
Effective IOH prediction requires capturing both short-term and long-term dependencies in MAP and SBP series. The Transformer encoder, with its self-attention mechanism, is well-suited for this task, dynamically weighting the importance of different time segments to model complex temporal patterns.
After obtaining patch embeddings $X_{\text{pos}}$ via patch embedding and positional encoding, the Transformer encoder processes them through self-attention and feedforward layers to learn sequence dependencies. Given positional embeddings $X_{\text{pos}}\in\mathbb{R}^{\frac{L}{S}\times d_{\text{model}}}$, the dependencies are computed as:
$Z=\text{TransformerEncoder}(W_{\text{pos}}),$ (8)
where $Z\in\mathbb{R}^{\frac{L}{S}\times d_{\text{model}}}$ encodes cross-segment relationships in MAP and SBP. This formulation ensures that the model captures subtle fluctuations and interactions essential for dynamic sequence modeling. By leveraging self-attention mechanism, the Transformer enhances predictive capability, aligning with the multi-factor feature learning framework to improve IOH forecasting.
Table 1: Statistics of training, validation, and testing sets across different datasets and prediction scopes.
Dataset Patient Number Sampling Rate (s) Predicted Scope Training Set Validation Set Testing Set
CH-OPBP 1,083 1 300 442,858 54,406 58,019
600 489,273 58,810 66,071
900 519,339 61,776 71,130
3 100 148,274 18,243 19,401
200 164,049 19,719 22,124
300 174,311 20,764 23,848
VitalDB 1,522 3 100 466,402 61,158 58,042
200 589,414 78,037 74,094
300 689,650 91,902 87,685
3.5 Prediction layer
To ensure that the temporal patterns captured by the Transformer encoder are effectively utilized in MAP series predictions, we employ a linear layer to map the sequence representation $Z$ to the predicted output. This transformation enables the model to preserve sequence dependencies while efficiently translating extracted features into precise forecasts. To maintain temporal coherence, we employ a recurrent forecasting strategy, where each predicted step conditions the next, ensuring consistency across the forecasted sequence. The mapping from $Z\in\mathbb{R}^{\frac{L}{S}\times d_{\text{model}}}$ to $Y\in\mathbb{R}^{\frac{L}{S}\times\frac{2TS}{L}}$ is formulated as:
$Y=Z\cdot W+b,$ (9)
where $W$ and $b$ denote the transformation matrix and bias vector, respectively. Linear transformations have been widely adopted in sequence forecasting Zeng et al. (2023), facilitating the direct conversion of encoded representations into accurate future MAP predictions.
3.6 Optimization Strategies
To enhance prediction accuracy, we incorporate a patch-based autoregressive method. The process begins with the final patch of the decomposed component $W$, denoted as $Y_{0}=W_{\frac{L}{S}}$, and sequentially generates subsequent patches using:
$P(Y^{\prime}_{i+1})=\prod_{j=1}^{i}P(Y^{\prime}_{j+1}|Y^{\prime}_{j}),$ (10)
in which $Y^{\prime}_{i+1}$ represents the next patch. The predicted sequence can be represented as:
$\hat{Y}^{\prime}=\{Y^{\prime}_{1},Y^{\prime}_{2},\dots,Y^{\prime}_{\frac{T}{S}}\}.$ (11)
By leveraging prior patches, this method improves temporal dependency modeling for MAP forecasting. The model is optimized by minimizing the mean squared error (MSE) between the predicted and actual sequences:
$\text{MSE}=\frac{1}{T}\sum_{t=1}^{T}\left(Y_{t}-\hat{Y}_{t}\right)^{2}.$ (12)
This optimization refines forecasting accuracy, enhancing IOH prediction effectiveness.
4 Experiments
4.1 Experimental Setup
Dataset Description.
We conduct experiments on two real-world datasets: CH-OPBP and VitalDB Lee (2018). The CH-OPBP dataset is collected from real-world intraoperative monitoring records of 3,422 patients, sampled at 100Hz. The data acquisition process strictly adheres to the ethical guidelines and protocols approved by the hospital ethics committee, ensuring compliance with patient privacy and regulatory standards. After resampling to 1-second and 3-second intervals and filtering out records shorter than one hour, 1,083 records remain. Data collection spans from February 27, 2023, to August 4, 2023. The VitalDB dataset initially includes 6,388 records, resampled at 3-second intervals. After excluding records with more than 20% missing data, 1,522 records remain. To ensure robust model evaluation, both datasets follow an 80%-10%-10% split for training, validation, and testing while maintaining temporal consistency. The CH-OPBP dataset supports multiple prediction horizons, whereas the VitalDB dataset focuses on 3-second sampled data. Table 1 summarizes data partitioning across different sampling rates and prediction scopes. Figure 4 visualizes the partitioning strategy, segmenting each patient’s surgical timeline for structured training and evaluation.
Figure 4: Visualization of data partitioning across different patients during surgery time.
Algorithm 1 IOH Events Detection
Input: $O_{\text{pred}}$: predicted sequence of MAP values, $t$: minimum duration of IOH, $T_{\text{instance}}$: length of the target window, $L_{\text{actual}}$: actual labels sequence.
Output: $J_{\text{actual}}$: boolean indicating the presence of an IOH event in the actual instance, $J_{\text{prediction}}$: boolean indicating the presence of an IOH event in the predicted instance.
1:  $J_{\text{actual}}=0$
2:  $J_{\text{prediction}}=0$
3:  for $i$ in range(0, $T_{\text{instance}}-t+1$do
4:     $\text{sum}_{\text{actual}}=\text{sum}(L_{\text{actual}}[i:i+t])$
5:     if $\text{sum}_{\text{actual}}>0$ then
6:        $J_{\text{actual}}=1$
7:     end if
8:     $\text{sum}_{\text{prediction}}=\sum_{j=0}^{t-1}(O_{\text{pred}}[i+j]\leq 65)$
9:     if $\text{sum}_{\text{prediction}}>0.8*t$ then
10:        $J_{\text{prediction}}=1$
11:     end if
12:  end for
13:  return $J_{\text{actual}}$, $J_{\text{prediction}}$
Table 2: Performance comparison of IOH prediction between our HMF and baseline models on two datasets .
Datasets Sample (s) Model MSE MAE AUC Accuracy (%) Recall (%)
CH-OPBP 1 Arima 130.2702 8.8526 0.5963 77.32 24.00
LR 0.5054 76.08 35.10
LSTM 118.1246 9.0985 0.5295 76.78 6.58
Transformer 126.7972 9.3809 0.5919 75.63 22.43
Informer 103.7028 8.0757 0.6452 72.26 35.38
DLinear 125.6786 9.3232 0.5331 71.60 7.54
HMF 93.2677 7.5823 0.7352 75.29 67.98
CH-OPBP 3 Arima 112.9281 8.1606 0.5928 75.70 25.65
LR 0.6774 75.71 54.49
LSTM 124.4213 9.8814 0.5000 75.53 0.00
Transformer 104.8545 8.3441 0.5970 74.44 23.12
Informer 111.0393 8.3883 0.6278 79.81 30.37
DLinear 123.8899 9.2951 0.5413 74.10 11.98
HMF 86.4927 7.2828 0.7413 61.38 70.13
VitalDB 3 Arima 257.3701 13.1127 0.5250 59.31 8.53
LR 0.5595 62.60 33.47
LSTM 188.7123 11.8613 0.5000 75.62 0.00
Transformer 158.7031 10.6901 0.5040 73.51 0.93
Informer 158.7873 10.8987 0.5003 75.01 0.05
DLinear 175.1144 11.4968 0.5074 65.09 1.86
HMF 165.7575 9.3845 0.6468 69.27 45.87
Table 3: Ablation study on instance normalization and sequence decomposition in HMF on the CH-OPBP dataset.
Dataset Model MSE MAE AUC Accuracy (%) Recall (%)
CH-OPBP HMF (full model) 86.4927 7.2828 0.7413 75.53 70.13
w/o instance normalization 106.9671 8.5005 0.5891 77.72 20.73
w/o sequence decomposition 105.7231 8.4964 0.5750 78.10 17.50
Compared Baselines.
To evaluate HMF, we compare it with traditional and deep learning-based forecasting models. For traditional methods, we use ARIMA Ariyo et al. (2014) and Logistic Regression. ARIMA, a statistical time series model, is trained on 0.5% of the test data for efficiency. Logistic Regression, a standard classifier, is enhanced with 1,566 features extracted via tsfresh [^4] https://github.com/blue-yonder/tsfresh and evaluated on 1% of the dataset. For deep learning baselines, we include LSTM Graves and Graves (2012), Transformer Vaswani et al. (2017), Informer Zhou et al. (2021), and DLinear Zeng et al. (2023). LSTM captures sequential dependencies via recurrence, while Transformer leverages self-attention for long-range interactions. Informer improves efficiency with sparse attention, making it suitable for long-sequence forecasting. DLinear, a lightweight alternative, models trend and seasonal components separately. Further implementation details and hyperparameter settings are in the supplementary material.
Implementation and Evaluation Details.
The HMF model predicts future MAP values based on physiological forecasts of both MAP and SBP, utilizing a 15-minute context window as input. Predictions extend over 5, 10, and 15 minutes to capture short- and long-term trends. Model training minimizes Mean Squared Error (MSE), with evaluation using both MSE and Mean Absolute Error (MAE), focusing on hypotensive segments ($L_{\text{actual}}=1$). Although both MAP and SBP series are predicted, IOH assessment relies solely on the forecasted MAP trajectory. Following clinical guidelines Wesselink et al. (2018), a hypotensive event is identified when predicted MAP remains below $\theta_{\text{MAP}}=65$ mmHg for at least 1 minute. To mitigate the impact of transient fluctuations, a 2-minute skip window Wijnberge et al. (2020) is applied, ensuring evaluation focuses on sustained hypotensive trends rather than isolated drops. Performance is assessed using accuracy, recall, and Area Under the Curve (AUC), with AUC serving as the primary metric for evaluating the model’s ability to differentiate hypotensive trends. Algorithm 1 details the IOH event detection process by comparing predicted and actual MAP sequences in a time-series forecasting framework. Further methodological details, including ground truth labeling and evaluation protocols, are provided in supplement material part.
4.2 Experimental Results
Main Results Analysis.
Table 2 presents a comprehensive evaluation of predictive models for dynamic IOH prediction, highlighting the advantages of the HMF model. The results indicate that HMF consistently outperforms baseline methods across datasets and sampling rates. Its ability to capture the intricate dynamics of blood pressure trends is reflected in improved predictive accuracy and classification performance. Notably, HMF demonstrates robustness in handling the non-stationary and complex nature of intraoperative blood pressure data, where traditional models often struggle. In the IOH prediction task, HMF shows strong performance in detecting hypotensive events. Its improvements across different sampling rates suggest adaptability to varying temporal granularities, a crucial factor in clinical applications. Compared to LSTM and Transformer, which perform well in some scenarios but face challenges with long sequence dependencies, HMF effectively models temporal dependencies and coupling effects. Informer and DLinear, despite their advantages in general forecasting tasks, show limitations in handling intraoperative data, particularly in recall performance. Overall, HMF maintains stable performance across diverse conditions, demonstrating its potential to enhance IOH prediction in complex clinical settings.
Ablation Study of Key Components.
To evaluate the impact of key components in the HMF framework, we conduct an ablation study on the CH-OPBP dataset. Table 3 presents the results, highlighting the roles of instance normalization and sequence decomposition in MAP forecasting and IOH detection. The full HMF model, integrating both components, achieves the best overall performance, suggesting their complementary effects in enhancing predictive capability. Removing instance normalization leads to a notable performance drop, indicating its importance in managing the non-stationary nature of intraoperative blood pressure data. Without normalization, the model struggles with variability, affecting its ability to detect IOH events consistently. Similarly, excluding sequence decomposition degrades performance, suggesting its role in capturing complex temporal dependencies. Without decomposition, the model may fail to identify underlying trends, limiting predictive accuracy. These findings underscore the importance of both components in improving model accuracy. Integrating instance normalization and sequence decomposition enhances generalization, leading to more effective IOH predictions.
Table 4: Comparison of transfer and non-transfer learning on individual patients.
Methods Patient ID MSE MAE AUC
Transfer 1 37.3299 5.6241 0.5375
2 39.9698 6.2213 0.6100
Non-Transfer 1 18.6823 3.8187 0.2375
2 47.6620 6.8317 0.5600
Table 5: Comparison of transfer and non-transfer learning on elderly and young subgroups.
Methods Feature MSE MAE AUC
Transfer Elderly 118.9462 9.1424 0.6089
Young 38.0568 5.0047 0.8040
Non-Transfer Elderly 66.5058 6.5874 0.7850
Young 44.1721 5.2447 0.8540
Performance Analysis w.r.t Transfer Settings.
Assessing model transferability across patients and demographic groups is essential for IOH prediction. Transfer learning enhances the robustness and generalizability of the HMF framework. To evaluate this, we apply the HMF model, trained on CH-OPBP, to new patient cohorts and age groups, comparing performance with and without transfer learning. Table 4 presents cross-patient transfer results, where a model trained on one patient is applied to another. AUC scores suggest that transfer learning preserves or even enhances predictive performance, indicating the model’s ability to capture shared physiological patterns. This reduces reliance on patient-specific data while maintaining accuracy, though individual training may still offer advantages in highly distinct cases. Table 5 presents cross-group transfer results between age cohorts. The non-transfer model performs better in elderly patients, suggesting that age-specific models better capture physiological complexities. In contrast, younger patients benefit from transfer learning, likely due to more homogeneous physiological responses. These findings underscore the importance of demographic-specific considerations while demonstrating transfer learning’s potential to improve predictive performance across diverse clinical settings.
Figure 5: Analysis of MSE and AUC performance for IOH segments across varying $d_{\text{model}}\in\{64,128,256,512,1024\}$.
Figure 6: Analysis of MSE and AUC performance for IOH segments across varying $S$$\in\{5,10,20,25,50\}$.
Parameter Sensitivity Analysis.
To optimize key model parameters and enhance the Patch Encoder’s ability to capture temporal dependencies, we conduct a parameter sensitivity analysis on the CH-OPBP dataset, focusing on the model dimension $d_{\text{model}}$ and patch length $S$. We first investigate the effect of varying $d_{\text{model}}$ while keeping the context window fixed at 300 and the predicted scope length at 100. As shown in Figure 5, different values of $d_{\text{model}}$ significantly influence model performance, with an optimal setting maximizing AUC. Subsequently, we examine the impact of varying $S$ while holding $d_{\text{model}}$ constant. This analysis is critical, as $S$ directly affects how effectively the model captures local temporal structures within physiological signals. The corresponding results, presented in Figure 6, indicate that selecting an appropriate $S$ value plays a crucial role in model performance. Based on these analyses, we determine the optimal parameter values that yield the highest AUC scores, ensuring robust model performance. The sensitivity of the model’s performance to $d_{\text{model}}$ and $S$ highlights the importance of appropriately configuring the Patch Encoder to accurately model blood pressure dynamics.
Visualization Case Study Analysis.
Figure 7 compares MAP predictions across models, demonstrating the effectiveness of the HMF framework. Our analysis of the CH-OPBP dataset, using a 15-minute context window and a 3-second sampling rate, consistently shows that HMF outperforms other models in predicting IOH events across various prediction horizons. Specifically, for a prediction length of 100, HMF (pink) closely follows actual MAP trends, capturing both rapid fluctuations and long-term variations. In contrast, DLinear (purple) and Informer (teal) tend to smooth out critical transitions, while Transformer (yellow) shows occasional delays. LSTM (orange) struggles with long-range dependencies, leading to deviations in extended forecasts. These results highlight HMF’s advantage in modeling complex temporal patterns for accurate IOH prediction.
Figure 7: Comparison of MAP prediction across different models on representative test cases.
5 Conclusion
We proposed a novel approach to IOH prediction by reformulating it as a time series forecasting task. To address key challenges, we introduced symmetric normalization to mitigate non-stationarity in MAP and SBP series, sequence decomposition to capture complex temporal patterns, and a patch-based Transformer model to enhance representation learning. Extensive experiments on real-world datasets validated the effectiveness of our HMF framework, demonstrating its ability to improve early IOH detection. This study highlights the potential of time series forecasting in clinical decision support and provides a foundation for future research in intraoperative risk prediction.
Appendix
Appendix A Related Work
A.1 Intraoperative Hypotension Forecasting
Existing methods for intraoperative hypotension (IOH) prediction predominantly adopt a static modeling approach, extracting handcrafted features from physiological signals rather than directly modeling their temporal dynamics. Early studies focused on high-fidelity arterial pressure waveforms, leading to the Hypotension Prediction Index (HPI) Hatib et al. [2018b]. Machine learning approaches, including ensemble methods Cherifa et al. [2020] and gradient boosting Kendale et al. [2018], integrated preoperative and intraoperative factors but largely treated data as independent samples, overlooking temporal dependencies. Deep learning models, such as RNN-based predictors Jeong et al. [2019] and attention-based frameworks Lu et al. [2023], improved performance by capturing sequential patterns but primarily relied on fixed-length feature representations, failing to fully model non-stationary hemodynamic changes. Recent interpretable models Ritter et al. [2023] enhanced clinical applicability but remained feature-driven. Our work addresses these limitations by formulating IOH prediction as a multivariate time-series forecasting problem, explicitly capturing temporal dependencies and non-stationary dynamics in arterial blood pressure signals.
A.2 Time Series and Sequence Forecasting
Time series forecasting is fundamental in domains such as finance, healthcare, and energy Cheng et al. [2024b]. Traditional models like ARIMA Ariyo et al. [2014] and exponential smoothing struggle with high-dimensional physiological signals He et al. [2023], while deep learning methods, including LSTMs Graves and Graves [2012] and GRUs Dey and Salem [2017], improve long-term dependency modeling. Recent Transformer-based models enhance long-sequence forecasting efficiency. Informer Zhou et al. [2021] optimizes self-attention mechanisms, Autoformer Chen et al. [2021] incorporates trend-seasonality decomposition, and FEDformer Zhou et al. [2022] leverages frequency domain analysis. Lightweight alternatives, such as DLinear Zeng et al. [2023] and TSMixer Ekambaram et al. [2023], demonstrate competitive performance by preserving temporal structures with simpler architectures. GPHT Liu et al. [2024] further introduces generative pretraining to improve forecasting transferability across datasets and horizons. Despite these advances Cheng et al. [2024b]; Zhang et al. [2024]; Cheng et al. [2024a]; Wang et al. [2024b, a], challenges remain in computational efficiency and real-time adaptability for physiological signal forecasting. Our work addresses these gaps by integrating multi-factor feature extraction with an efficient Transformer-based approach, improving IOH prediction accuracy in dynamic clinical settings.
Appendix B Evaluation Details
B.1 Ground Truth Labeling for IOH Events
In this study, ground truth labels for intraoperative hypotension (IOH) events were generated based on recorded mean arterial pressure (MAP) values during surgery. The labeling process identifies clinically significant hypotensive episodes while minimizing transient fluctuations that may not reflect sustained hemodynamic instability.
An IOH event is defined as a segment where the MAP value remains below a predefined threshold $\theta_{\text{MAP}}$ for at least $t$ minutes. Following clinical guidelines Wesselink et al. [2018], we set $\theta_{\text{MAP}}=65$ mmHg, a widely accepted hypotension threshold, and $t=1$ minute to ensure clinical relevance.
To generate labels, the algorithm scans the MAP time series for each patient, evaluating non-overlapping segments of length $t$. For each segment, the maximum MAP value is determined. If this value is below $\theta_{\text{MAP}}$, the entire segment is labeled as an IOH event. This approach ensures that only sustained hypotensive episodes are captured, reducing false positives from transient MAP drops. The labeling process is outlined in Algorithm 2.
Algorithm 2 Ground Truth Labeling for IOH Events
Input: $MAP_{\text{actual}}$: recorded MAP series, $t$: minimum duration for an IOH event, $T_{\text{seq}}$: total length of the MAP series.
Output: $L_{\text{actual}}$: binary label sequence for IOH events.
1:  Initialize $L_{\text{actual}}[:]=0$
2:  for $i$ in range(0, $T_{\text{seq}}-t+1$do
3:     $MAP_{\text{max}}=\max(MAP_{\text{actual}}[i:i+t])$
4:     if $MAP_{\text{max}}\leq\theta_{\text{MAP}}$ then
5:        $L_{\text{actual}}[i:i+t]=1$
6:     end if
7:  end for
8:  return $L_{\text{actual}}$
B.2 Evaluation of Hypotensive Events
To evaluate IOH detection, we compare the model’s predicted MAP trajectory against actual MAP values, assessing both event identification and duration. A predicted IOH event is correct if the forecasted MAP remains below $\theta_{\text{MAP}}$ for at least $t$ minutes.
Performance is quantified using the following metrics, including precision, recall, and F1-score:
$\text{Precision}=\frac{\text{TP}}{\text{TP}+\text{FP}}$ (13)
$\text{Recall}=\frac{\text{TP}}{\text{TP}+\text{FN}}$ (14)
[…]
where TP (True Positives) are correctly identified IOH events, FP (False Positives) are incorrect predictions, and FN (False Negatives) are missed hypotensive episodes.
Additionally, we compute the time-to-detection (TTD), which measures how early the model predicts an IOH event before its actual onset. A lower TTD indicates better early-warning capability, which is essential for timely clinical intervention.
This evaluation framework ensures that our method provides meaningful predictions, distinguishing sustained hypotensive episodes from transient MAP fluctuations while maintaining high predictive accuracy.
Appendix C Details of Baseline Methods
To comprehensively evaluate the performance of HMF in early IOH prediction, we benchmarked it against a set of traditional and deep learning-based methods. These baselines were selected to cover a range of modeling approaches, from statistical methods to advanced sequence learning techniques.
C.1 Traditional Methods
We included two widely used traditional time series forecasting and classification models:
  • ARIMA Ariyo et al. [2014]: The Autoregressive Integrated Moving Average (ARIMA) model was used as a classical statistical forecasting baseline. Due to its computational intensity, we trained ARIMA on only 0.5% of the test set, balancing efficiency and representativeness in the comparison.
  • Logistic Regression: A standard binary classification model, logistic regression was applied to predict IOH events. To enhance predictive performance, time series features were extracted using the tsfresh library [^5] https://github.com/blue-yonder/tsfresh, which generated 1,566 handcrafted features, capturing statistical and frequency-domain properties.
C.2 Deep Learning Methods
We evaluated several deep learning-based baselines, each known for its effectiveness in modeling sequential dependencies in time series data:
  • LSTM Graves and Graves [2012]: The Long Short-Term Memory (LSTM) network, a recurrent neural network (RNN) variant, was included due to its capability to capture long-range temporal dependencies. Its gating mechanism mitigates vanishing gradient issues, making it suitable for sequential modeling tasks like IOH prediction.
  • GRU Dey and Salem [2017]: The Gated Recurrent Unit (GRU) is an RNN variant similar to LSTM but with a simplified architecture. GRU was included for comparison due to its efficiency and comparable performance in capturing time series patterns while using fewer parameters.
  • Transformer Vaswani et al. [2017]: The Transformer model, with its self-attention mechanism, was included to evaluate the effectiveness of attention-based architectures in IOH prediction. By allowing the model to weigh different time steps adaptively, it can capture complex dependencies beyond what RNN-based models achieve.
  • Informer Zhou et al. [2021]: The Informer model extends the Transformer architecture by introducing a ProbSparse self-attention mechanism, reducing computational cost while preserving long-range dependency modeling. Its efficiency makes it well-suited for long-sequence forecasting tasks.
  • DLinear Zeng et al. [2023]: A lightweight linear model for time series forecasting, DLinear models trend and seasonal components separately. Despite its simplicity, it has demonstrated competitive performance against complex deep learning models, providing a strong baseline.
  • TSMixer Ekambaram et al. [2023]: A recent MLP-based model, TSMixer improves time series forecasting by using a structured mixing approach for temporal dependencies. We included it to assess the performance of purely feedforward architectures compared to sequential and attention-based models.
These baselines provide a diverse comparison set, ranging from traditional statistical methods to state-of-the-art deep learning architectures. By including both simple and complex models, we ensure a rigorous evaluation of HMF’s effectiveness in IOH prediction.
Table 6: Performance Comparison of Different Methods Using The CH-OPBP Dataset.
Model Sample(s) Pred MSE MAE AUC Accuracy (%) Recall (%)
Arima 1 300 97.1580 7.5501 0.6071 90.34 23.33
600 118.0238 8.8539 0.6022 75.15 25.53
900 175.6289 10.1539 0.5795 66.48 23.13
3 100 26.9210 4.0537 0.6625 86.60 38.46
200 155.1008 10.0347 0.5870 77.27 22.22
300 156.7626 10.3933 0.5288 63.03 16.28
LR 1 300 0.5000 86.06 0.00
600 0.6777 62.13 81.90
900 0.3386 35.96 23.39
3 100 0.5977 79.84 32.50
200 0.7002 72.00 65.18
300 0.7342 75.01 65.79
LSTM 1 300 98.2363 7.9263 0.5660 90.65 14.57
600 122.3699 9.4033 0.5000 73.49 0.00
900 133.7677 9.9659 0.5226 64.47 5.18
3 100 104.2937 8.9187 0.5000 90.51 0.00
200 127.8037 10.0756 0.5000 73.48 0.00
300 141.1666 10.6498 0.5000 62.94 0.00
Transformer 1 300 106.0975 8.2478 0.6194 89.99 27.32
600 126.0585 9.6861 0.5804 75.94 19.94
900 148.2355 10.2089 0.5758 67.32 20.04
3 100 93.5870 7.4800 0.6328 89.51 30.91
200 112.6674 9.0762 0.5406 75.06 9.34
300 108.3092 8.4761 0.6177 70.22 29.11
Informer 1 300 102.4836 8.2318 0.5663 90.41 14.95
600 92.3734 7.6706 0.6952 79.20 48.91
900 116.2513 8.4248 0.6741 73.93 42.27
3 100 104.9414 8.1057 0.6065 90.11 24.28
200 108.3698 8.7226 0.5934 77.23 21.25
300 119.7905 8.3365 0.6835 74.24 45.59
DLinear 1 300 107.2025 8.2388 0.5468 90.39 10.61
600 127.4571 9.6158 0.5316 74.75 7.20
900 142.3762 10.1150 0.5208 64.33 4.82
3 100 106.1303 8.2176 0.5587 90.27 13.42
200 123.7473 9.4676 0.5711 76.31 16.21
300 141.7650 10.2002 0.5271 64.72 6.30
HMF 1 300 73.7196 6.8555 0.7702 76.48 77.70
600 89.8372 7.6161 0.7230 75.30 65.91
900 116.2464 8.2751 0.7125 74.08 60.33
3 100 77.0390 6.8708 0.7801 77.72 78.38
200 88.6844 7.4293 0.7245 75.32 66.34
300 93.7548 7.5482 0.7194 73.55 65.68
Table 7: Performance Comparison of Different Methods Using The VitalDB Dataset.
Model Sample(s) Pred MSE MAE AUC Accuracy(%) Recall(%)
Arima 3 100 193.3889 10.7640 0.5397 83.10 10.42
200 268.3629 13.7636 0.5073 60.27 5.48
300 310.3584 14.8105 0.5280 53.20 9.68
LR 3 100 0.5094 78.52 2.11
200 0.5877 61.80 37.50
300 0.5814 58.41 60.80
LSTM 3 100 157.8527 10.6642 0.5000 84.35 0.00
200 204.0663 12.3266 0.5000 61.35 0.00
300 204.2178 12.5930 0.5000 48.58 0.00
Transformer 3 100 162.8149 10.6928 0.5022 84.36 0.54
200 162.6087 10.8778 0.5006 61.40 0.13
300 150.6858 10.4997 0.5091 49.50 2.12
Informer 3 100 164.9233 11.0281 0.5000 84.35 0.00
200 152.0211 10.7310 0.5000 61.35 0.00
300 159.4176 10.9371 0.5008 48.66 0.16
DLinear 3 100 173.4594 11.1419 0.5097 84.33 2.41
200 175.6332 11.4898 0.5077 61.85 1.97
300 176.2505 11.8588 0.5049 49.08 1.19
HMF 3 100 149.5450 8.7662 0.6819 78.29 53.50
200 165.2017 9.3708 0.6370 68.01 44.74
300 182.5259 10.0166 0.6215 61.50 39.38
Appendix D Appendix D: Full Results Report
Tables 6 and 7 present the full experimental results comparing the performance of different baseline methods and our proposed HMF model on the CH-OPBP and VitalDB datasets. These tables provide a detailed evaluation across multiple metrics, including mean squared error (MSE), mean absolute error (MAE), area under the curve (AUC), accuracy, and recall, under different prediction horizons (e.g., 100, 200, and 300 sample steps). The results highlight the effectiveness of HMF in capturing intraoperative hypotension (IOH) dynamics compared to traditional and deep learning-based approaches. Notably, HMF demonstrates superior performance in key metrics such as AUC and recall, indicating its ability to better identify hypotensive events while maintaining predictive accuracy.

### v4

@@
 Abstract
-Intraoperative hypotension (IOH) prediction using past physiological signals is crucial, as IOH can lead to inadequate organ perfusion, increasing the risk of severe complications and mortality.
-However, existing IOH prediction methods often rely on static modeling, overlooking the complex temporal dependencies and non-stationary nature of physiological signals.
-In this paper, we propose a Hybrid Multi-Factor (HMF) network that models IOH prediction as a dynamic sequence forecasting problem, explicitly capturing temporal dependencies and physiological non-stationarity..
-Specifically, we formalize physiological signal dynamics as a sequence of multivariate time series, and decompose them into trend and seasonal components, enabling distinct modeling of long-term and periodic variations.
-For each component, we employ a patch-based Transformer encoder to extract representative features with the concern of computational efficiency and representation quality.
-Furthermore, to mitigate distributional drift arising from the evolving signals, we introduce a symmetric normalization mechanism.
-Extensive experiments on both a publicly available dataset and a private dataset collected from real-world hospital settings demonstrate that our approach significantly outperforms competitive baselines.
-We hope HMF offers a new perspective on IOH prediction and further enhances surgical safety [^2] https://github.com/Mingyue-Cheng/HMF.
+Intraoperative hypotension (IOH) prediction using past physiological signals is crucial, as IOH may lead to inadequate organ perfusion and significantly elevate the risk of severe complications and mortality.
+However, current methods often rely on static modeling, overlooking the complex temporal dependencies and the inherently non-stationary nature of physiological signals.
+We propose a Hybrid Multi-Factor (HMF) network that formulates IOH prediction as a dynamic sequence forecasting task, explicitly capturing both temporal dependencies and physiological non-stationarity.
+We represent signal dynamics as multivariate time series and decompose them into trend and seasonal components, enabling separate modeling of long-term and periodic variations.
+Each component is encoded with a patch-based Transformer to balance computational efficiency and feature representation.
+To address distributional drift from evolving signals, we introduce a symmetric normalization mechanism.
+Experiments on both public and real-world clinical datasets show that HMF significantly outperforms competitive baselines.
+We hope HMF offers new insights into IOH prediction and ultimately promotes safer surgical care.
+Our code is available at https://github.com/Mingyue-Cheng/HMF.
 1 Introduction
-Figure 1: Illustration of MAP Sequence: Temporal Dynamics and Distribution Shifts in Intraoperative Hypotension Event.
+Figure 1: Illustration of MAP sequence: temporal dynamics and distribution shifts in intraoperative hypotension events.
 Intraoperative mortality has decreased by a factor of $100$ over the past century, making deaths during surgery a rare occurrence Li et al.
@@
 (2019)—amounting to more than $4$ million deaths worldwide each year Saugel and Sessler (2021).
-These postoperative deaths are most strongly linked to complications, which are often triggered by intraoperative events.
-Among these, intraoperative hypotension (IOH) is a common and serious complication Kim et al.
-(2023), characterized by a significant drop in Mean Arterial Pressure (MAP) [^3] MAP, derived from Arterial Blood Pressure (ABP), reflects tissue perfusion and is calculated as $MAP=DBP+\frac{1}{3}(SBP-DBP)$, where SBP and DBP denote peak and minimum arterial pressures per cardiac cycle.
-A sustained decline in SBP and DBP reduces MAP, risking organ hypoperfusion.
-sustained over a period of time.
+These postoperative deaths are most strongly linked to complications Lee et al.
+(2017), which are often triggered by intraoperative events.
+Among these, intraoperative hypotension (IOH) is a prevalent and clinically significant complication Kim et al.
+(2023), characterized by a sustained reduction in mean arterial pressure  [^2] MAP, derived from arterial blood pressure (ABP), is a key indicator of tissue perfusion.
+It is calculated as $MAP=DBP+\frac{1}{3}(SBP-DBP)$ Gauer (1960), where systolic blood pressure (SBP) and diastolic blood pressure (DBP) represent the maximum and minimum arterial pressures within a cardiac cycle, respectively.
+(MAP).
 Previous studies Hatib et al.
 (2018a) have shown that IOH events can be predicted using machine learning applied to physiological signals.
-Clinical trials further indicate that early warnings and timely interventions help mitigate hypotension severity and reduce postoperative complications Hwang et al.
+Clinical trials further indicate that early warnings and timely interventions help reduce hypotension severity and postoperative complications Hwang et al.
 (2023); Fernandes et al.
 (2021).
-Recent advances in IOH prediction primarily follow a feature extraction-based approach Lee et al.
+Recent advances primarily rely on handcrafted features combined with classifiers such as logistic regression and random forests Lee et al.
 (2021); Davies et al.
-(2020), where handcrafted features are used with classifiers like logistic regression and random forests.
-However, these methods face two key limitations: (1) reliance on handcrafted features limits their ability to capture temporal dependencies, and (2) their models lack the flexibility to adapt to complex real-world IOH scenarios.
-Additionally, IOH definitions vary across clinical settings, and fixed threshold-based criteria may fail to accurately characterize IOH, particularly in hypertensive or elderly patients, where more nuanced definitions are often needed.
-To overcome the limitations of existing approaches, we reformulate IOH prediction as a multivariate time series forecasting problem.
-The core rationale behind this formulation is that time series modeling effectively captures the temporal dynamics and distribution shifts inherent in physiological signals.
-Since hypotensive events are defined based on MAP series, accurate sequence forecasting directly impacts risk identification, offering greater flexibility.
-However, as illustrated in Figure 1, several challenges arise in MAP series prediction.
-First, the temporal dynamic waveform of MAP consists of multiple components, including trends and periodic patterns, which may need to be modeled separately to improve forecasting accuracy.
-Second, the high information redundancy caused by the extended length of MAP waveforms increases computational complexity and complicates representation learning in sequence modeling.
-Finally, evolving distribution shifts introduce additional challenges, as sudden fluctuations in blood pressure alter the statistical properties (e.g., mean and standard deviation) of the MAP series, making it difficult to maintain consistent predictive performance.
-Addressing these challenges is essential for developing a high-accurate and adaptable IOH prediction framework.
-To address these limitations, we introduce the Hybrid Multi-Factor (HMF) framework, which explicitly models temporal dependencies and mitigates distribution shifts, ensuring adaptive and accurate IOH prediction.
-A key feature of HMF is its use of sequence decomposition techniques to model the trend and seasonal components separately, allowing for a more structured representation of historical dynamics.
-By integrating multi-factor physiological signals, including MAP and SBP, HMF effectively captures both local fluctuations and long-term dependencies.
-To efficiently extract informative features, we adopt a patch-based Transformer encoder to learn representations of the given sequence.
-Additionally, a symmetric normalization mechanism is introduced to mitigate distribution shifts caused by evolving physiological signals.
-We conduct extensive experiments on both public and private datasets collected from real-world clinical scenarios.
-The results demonstrate the superiority of HMF in improving predictive performance while effectively modeling multi-factor temporal interactions.
+(2020), but face two key limitations: (1) limited ability to capture temporal dependencies, and (2) lack of flexibility to handle complex real-world IOH scenarios.
+Additionally, fixed threshold-based IOH definitions may fail to reflect risk accurately in hypertensive or elderly patients, where more nuanced criteria are often needed.
+To overcome these limitations, we reformulate IOH prediction as a multivariate time series forecasting problem.
+This formulation captures the temporal dynamics and evolving distributions inherent in physiological signals.
+Since hypotensive events are defined based on MAP series, accurate forecasting directly enhances risk identification.
+However, as illustrated in Figure 1, several challenges arise.
+First, MAP waveforms comprise trends and periodic components, which may require separate modeling for improved accuracy.
+Second, redundancy in long MAP sequences increases computational cost and complicates representation learning.
+Third, sudden fluctuations alter statistical properties such as mean and standard deviation, making it difficult to maintain consistent performance.
+Addressing these challenges is essential for a robust and adaptive IOH prediction framework.
+We propose the Hybrid Multi-Factor (HMF) framework, which explicitly models temporal structure and mitigates distribution shifts.
+HMF employs sequence decomposition to separate trend and seasonal components, yielding a more structured representation of historical dynamics.
+By incorporating multi-factor physiological signals such as MAP and SBP, HMF captures both short-term fluctuations and long-range dependencies.
+A patch-based Transformer encoder is adopted for efficient sequence representation, while a symmetric normalization mechanism compensates for evolving input distributions.
+Extensive experiments on public and private intraoperative datasets demonstrate that HMF consistently improves predictive performance and effectively models multi-factor temporal interactions.
 The key contributions are summarized as follows:
@@
   • We validate the effectiveness of the proposed approach through extensive experiments on two real-world datasets, demonstrating superior performance compared to existing methods.
+2 Related Work
+2.1 Intraoperative Hypotension Forecasting
+Existing methods for intraoperative hypotension (IOH) prediction predominantly adopt a static modeling approach, extracting handcrafted features from physiological signals rather than directly modeling their temporal dynamics.
+Early studies focused on high-fidelity arterial pressure waveforms, leading to the Hypotension Prediction Index (HPI) Hatib et al.
+(2018b).
+Machine learning approaches, including ensemble methods Cherifa et al.
+(2020) and gradient boosting Kendale et al.
+(2018), integrated preoperative and intraoperative factors but largely treated data as independent samples, overlooking temporal dependencies.
+Deep learning models, such as RNN-based predictors Jeong et al.
+(2019) and attention-based frameworks Lu et al.
+(2023), improved performance by capturing sequential patterns but primarily relied on fixed-length feature representations, failing to fully model non-stationary hemodynamic changes.
+Recent interpretable models Ritter et al.
+(2023) enhanced clinical applicability but remained feature-driven.
+Our work addresses these limitations by formulating IOH prediction as a multivariate time-series forecasting problem, explicitly capturing temporal dependencies and non-stationary dynamics in arterial blood pressure signals.
+2.2 Time Series and Sequence Forecasting
+Time series forecasting is fundamental in domains such as finance, healthcare, and energy Cheng et al.
+(2025, 2024).
+Traditional models like ARIMA Hyndman and Khandakar (2008) and exponential smoothing struggle with high-dimensional physiological signals, while deep learning methods, including LSTMs Graves and Graves (2012) and GRUs Dey and Salem (2017), improve long-term dependency modeling.
+Recent Transformer-based models enhance long-sequence forecasting effectiveness.
+Informer Zhou et al.
+(2021) optimizes self-attention mechanisms through sparse attention and Autoformer Chen et al.
+(2021) leverages series decomposition to explicitly model trend and seasonal components, which enhances interpretability and improves forecasting stability over long horizons.
+Lightweight alternatives, such as DLinear Zeng et al.
+(2023), demonstrate competitive performance by preserving temporal structures with simpler architectures.
+Due to the data sparsity issue Cheng et al.
+(2023), GPHT Liu et al.
+(2024) further introduces generative pretraining to improve forecasting transferability across datasets and horizons.
+While significant progress has been made in deterministic forecasting, probabilistic approaches such as diffusion-based models Tashiro et al.
+(2021) Zhang et al.
+(2025) are also gaining increasing attention for their ability to capture uncertainty.
+Complementarily, supervised architectures such as TimeDART Wang et al.
+(2025) highlight the potential of discriminative learning in capturing temporal patterns effectively.
+On the other hand, the introduction of LLM-based models Jin et al.
+(2023) Zhou et al.
+(2023) has enabled the integration of multimodal information.
+These advancements collectively provide a solid foundation for more accurate modeling of complex medical time series.
+While existing models often struggle with real-time adaptability and the challenges posed by non-stationary physiological sequences in dynamic clinical settings, our work bridges this gap by combining efficient patch-based sequence encoding with structured modeling of physiological signals.
+This design enables rapid inference and robust adaptation to evolving patient states, thereby improving IOH prediction accuracy in dynamic clinical settings.
 Figure 2: Feature extraction and temporal window setup for blood pressure trend prediction: context, skipped, and target windows.
-Figure 3: Proposed Hybrid Multi-Factor Model Architecture for Intraoperative Hypotension Prediction.
+Figure 3: Proposed hybrid multi-factor model for intraoperative hypotension prediction.
 3 Preliminaries
-To formulate the IOH prediction task, we first define how physiological features are extracted and structured, which serves as the foundation for subsequent problem formulation.
-2.1 Multi-factor Feature Construction
-To predict intraoperative hypotension (IOH), real-time monitoring and modeling techniques analyze arterial blood pressure (ABP) dynamics.
-Identifying early patterns preceding hypotensive events is crucial for timely intervention.
-This study structures multi-factor ABP-derived features to enhance forecasting accuracy.
-During surgery, ABP is continuously monitored at high frequency (e.g., 100 Hz), capturing hemodynamic fluctuations.
-Prior studies reveal a strong correlation between ABP variations within a context window and subsequent hypotension.
-Clinically, hypotension is defined based on Mean Arterial Pressure (MAP), the average pressure per cardiac cycle, considered hypotensive when sustained below a threshold (e.g., one minute).
-Systolic Blood Pressure (SBP), the peak pressure per cycle, is included due to its physiological link to MAP.
-These ABP-derived signals—MAP and SBP—serve as core predictive features.
-To construct informative representations, we extract continuous ABP segments, removing outliers and artifacts.
-A sliding window approach processes ABP series, extracting SBP (peak pressure) and Diastolic Blood Pressure (DBP, lowest pressure).
-MAP is computed as:
-$\text{MAP}=\frac{1}{3}\times\text{SBP}+\frac{2}{3}\times\text{DBP}.$ (1)
-Hypotensive events are identified when MAP remains below 65 mmHg for at least one minute, following clinical guidelines.
-Event annotation details, including threshold selection and temporal alignment, are further provided in the experimental and supplementary material.
-For IOH prediction, we construct a multivariate time series dataset incorporating SBP and MAP to capture both short-term fluctuations and long-term trends.
-This formulation enables effective feature extraction and temporal alignment, as illustrated in Figure 2.
-By integrating multiple ABP-derived factors, our approach improves the model’s ability to learn temporal dependencies and physiological dynamics associated with IOH.
+3.1 Multi-factor Sequence Construction
+To predict intraoperative hypotension (IOH), we analyze arterial blood pressure (ABP) dynamics to identify early indicators that precede hypotensive events, thereby enabling timely clinical intervention and risk mitigation.
+During surgery, ABP is continuously monitored at high frequency, capturing critical hemodynamic fluctuations.
+Prior studies have demonstrated a strong correlation between short-term variations in ABP and the onset of IOH.
+Clinically, IOH is defined as a period when the mean arterial pressure (MAP) remains below 65 mmHg for at least one minute.
+Systolic blood pressure (SBP), the peak value in a cardiac cycle, is also considered due to its physiological relation to MAP.
+These two signals, MAP and SBP, serve as the primary predictive inputs for IOH forecasting.
+To construct reliable input sequences we extract clean ABP segments using a sliding window, remove artifacts and compute SBP and diastolic blood pressure (DBP), as shown in the Figure 2.
 3.2 Problem Definition
-Building on the multi-factor feature construction framework, we formulate IOH risk prediction as a dynamic sequence forecasting problem.
-Given real-time arterial blood pressure (ABP) monitoring, our goal is to predict future mean arterial pressure (MAP) trends, enabling early warnings for potential hypotensive events.
-Formally, let $\mathbf{X}_{\text{MAP}}=\{x_{\text{MAP}}^{1},x_{\text{MAP}}^{2},\dots,x_{\text{MAP}}^{T}\}$ and $\mathbf{X}_{\text{SBP}}=\{x_{\text{SBP}}^{1},x_{\text{SBP}}^{2},\dots,x_{\text{SBP}}^{T}\}$ denote the historical MAP and systolic blood pressure (SBP) sequences over a context window of length $T$.
-The objective is to forecast the future MAP values $\mathbf{\hat{Y}}_{\text{MAP}}=\{\hat{y}_{\text{MAP}}^{T+1},\dots,\hat{y}_{\text{MAP}}^{T+\tau}\}$ over a prediction horizon $\tau$.
-A hypotensive event occurs if the predicted MAP falls below a predefined threshold $\theta_{\text{MAP}}$.
-Further details on event annotation and threshold selection are provided in supplement material part.
-By leveraging multi-factor temporal dependencies, our formulation enhances the reliability of IOH risk assessment, facilitating timely clinical intervention.
+Building on the multi-factor feature construction framework, we formulate IOH risk prediction as a dynamic sequence forecasting task.
+Given real-time ABP monitoring, the goal is to forecast future mean arterial pressure (MAP) trends to enable early warnings of hypotensive events.
+Let $\mathbf{X}_{\text{MAP}}=\{x_{\text{MAP}}^{1},\dots,x_{\text{MAP}}^{T}\}$ and $\mathbf{X}_{\text{SBP}}=\{x_{\text{SBP}}^{1},\dots,x_{\text{SBP}}^{T}\}$ represent historical MAP and systolic blood pressure (SBP) sequences over a context window of length $T$.
+The objective is to predict MAP values $\mathbf{\hat{Y}}_{\text{MAP}}=\{\hat{y}_{\text{MAP}}^{T+1},\dots,\hat{y}_{\text{MAP}}^{T+\tau}\}$ over horizon $\tau$.
+A hypotensive event is flagged if predicted MAP drops below a threshold $\theta_{\text{MAP}}$.
+Further details on annotation and thresholding are provided in the supplementary material.
+By modeling multi-factor temporal dependencies, our approach improves IOH risk assessment and supports timely clinical decisions.
 4 The Proposed HMF
-This section provides a detailed description of the proposed HMF framework.
 4.1 Framework Overview
@@
 where $\mu_{X}$ and $\sigma_{X}$ are the mean and standard deviation of $X$.
-This transformation allows the model to focus on relative changes rather than absolute values, reducing the impact of abrupt fluctuations and improving pattern recognition.
+This transformation allows the model to focus on relative changes rather than absolute values, reducing the impact of abrupt fluctuations.
 Additionally, by normalizing across instances, the method mitigates inter-patient variability, ensuring stable feature distributions.
@@
 This ensures stability while preserving raw characteristics, facilitating reliable and interpretable IOH forecasting.
-3.3 Sequence Decomposition Preprocessing
+4.3 Decomposition Layer
 The Arterial Blood Pressure(ABP) signal exhibits a complex waveform structure, consisting of multiple components like trend and periodic elements, which need to be individually modeled for accurate forecasting.
@@
 where AvgPool smooths the sequence by downsampling through averaging within a specified window, while Padding ensures full coverage across the entire sequence.
-3.4 Dynamic Feature Representation Layer
+4.4 Representation Layer
 Patch Embedding.
-To model the intricate dynamics of MAP and SBP series during surgery, we apply a patch embedding technique after the decomposition layer.
+To model the intricate dynamics of MAP and SBP series during surgery, we apply a patch embedding technique Nie et al.
+(2022) after the decomposition layer.
 This reduces sequence length, lowering computational complexity and improving efficiency.
@@
 The decomposed series $W\in\mathbb{R}^{L\times 2}$ is transformed into a compact representation $W_{\text{patch}}\in\mathbb{R}^{\frac{L}{S}\times d_{\text{model}}}$ using three 1D convolutional layers:
-$W_{\text{patch}}=\text{Conv1D}_{3}(W),$ (6)
-[…]
+$W_{\text{patch}}=\text{Conv1D}(W),$ (5)
+where Conv1D applies three consecutive 1D convolutions to $W$. The embedding dimension $d_{\text{model}}$ is chosen to extract key features, improving IOH prediction.
 To encode temporal structures, we introduce learnable positional encodings:
@@
 Sequence Dependence Modeling.
+Capturing cross channel dependencies is essential for accurate prediction of multivariate physiological time series Han et al.
+(2024).
 Effective IOH prediction requires capturing both short-term and long-term dependencies in MAP and SBP series.
@@
 By leveraging self-attention mechanism, the Transformer enhances predictive capability, aligning with the multi-factor feature learning framework to improve IOH forecasting.
-Table 1: Statistics of training, validation, and testing sets across different datasets and prediction scopes.
 Dataset Patient Number Sampling Rate (s) Predicted Scope Training Set Validation Set Testing Set
@@
 300 689,650 91,902 87,685
+Table 1: Statistics of training, validation, and testing sets across different datasets and prediction scopes.
 4.5 Prediction Layer
-To ensure that the temporal patterns captured by the Transformer encoder are effectively utilized in MAP series predictions, we employ a linear layer to map the sequence representation $Z$ to the predicted output.
-This transformation enables the model to preserve sequence dependencies while efficiently translating extracted features into precise forecasts.
-To maintain temporal coherence, we employ a recurrent forecasting strategy, where each predicted step conditions the next, ensuring consistency across the forecasted sequence.
+To ensure the temporal patterns captured by the Transformer encoder are effectively utilized in MAP predictions, we employ a linear layer to map the sequence representation $Z$ to the predicted output.
+This transformation allows the model to preserve sequence dependencies while translating extracted features into precise forecasts.
+To maintain temporal coherence, we use a recurrent forecasting strategy where each predicted step conditions the next, ensuring consistency across the forecasted sequence.
 The mapping from $Z\in\mathbb{R}^{\frac{L}{S}\times d_{\text{model}}}$ to $Y\in\mathbb{R}^{\frac{L}{S}\times\frac{2TS}{L}}$ is formulated as:
@@
 4.6 Optimization Strategies
-To enhance prediction accuracy, we incorporate a patch-based autoregressive method.
-The process begins with the final patch of the decomposed component $W$, denoted as $Y_{0}=W_{\frac{L}{S}}$, and sequentially generates subsequent patches using:
+To improve prediction accuracy, we integrate a patch-based autoregressive method into our model.
+This approach begins with the final patch of decomposed component $W$, denoted as $Y_{0}=W_{\frac{L}{S}}$, and sequentially generates subsequent patches using the autoregressive formula:
 $P(Y^{\prime}_{i+1})=\prod_{j=1}^{i}P(Y^{\prime}_{j+1}|Y^{\prime}_{j}),$ (9)
@@
 This optimization refines forecasting accuracy, enhancing IOH prediction effectiveness.
+Figure 4: Visualization of data partitioning across different patients during surgery time.
 5 Experiments
@@
 Dataset Description.
-We conduct experiments on two real-world datasets: CH-OPBP and VitalDB Lee (2018).
-The CH-OPBP dataset is collected from real-world intraoperative monitoring records of 3,422 patients, sampled at 100Hz.
-The data acquisition process strictly adheres to the ethical guidelines and protocols approved by the hospital ethics committee, ensuring compliance with patient privacy and regulatory standards.
-After resampling to 1-second and 3-second intervals and filtering out records shorter than one hour, 1,083 records remain.
-Data collection spans from February 27, 2023, to August 4, 2023.
-The VitalDB dataset initially includes 6,388 records, resampled at 3-second intervals.
-After excluding records with more than 20% missing data, 1,522 records remain.
-To ensure robust model evaluation, both datasets follow an 80%-10%-10% split for training, validation, and testing while maintaining temporal consistency.
-The CH-OPBP dataset supports multiple prediction horizons, whereas the VitalDB dataset focuses on 3-second sampled data.
+We conduct experiments on two real-world clinical datasets: CH-OPBP and VitalDB Lee (2018).
+The CH-OPBP dataset is collected from real-world intraoperative monitoring records of 3,422 patients, sampled continuously at 100Hz.
+The data acquisition process strictly adheres to ethical guidelines and protocols approved by the hospital ethics committee, ensuring full compliance with patient privacy regulations and relevant clinical research standards.
+After resampling to 1-second and 3-second intervals and filtering out records shorter than one hour, a total of 1,083 valid records remain.
+The data acquisition process also strictly follows patient de-identification protocols and has been formally approved by the Institutional Review Board of Anhui Provincial Hospital.
+Data collection spans the period from February 27, 2023, to August 4, 2023.
+The VitalDB dataset initially includes 6,388 intraoperative records, resampled uniformly at 3-second intervals.
+After excluding records with more than 20% missing data, 1,522 records are retained for use.
+To ensure reliable model evaluation, both datasets adopt an 80%-10%-10% split for training, validation, and testing, while preserving temporal consistency.
+The CH-OPBP dataset supports multiple prediction horizons to evaluate forecasting flexibility, whereas the VitalDB dataset focuses specifically on 3-second resolution data.
 Table 1 summarizes data partitioning across different sampling rates and prediction scopes.
-Figure 4 visualizes the partitioning strategy, segmenting each patient’s surgical timeline for structured training and evaluation.
-Figure 4: Visualization of data partitioning across different patients during surgery time.
+Figure 4 visualizes the partitioning strategy, segmenting each patient’s surgical timeline for structured model training and evaluation.
 Algorithm 1 IOH Events Detection
@@
 13:  return $J_{\text{actual}}$, $J_{\text{prediction}}$
-Table 2: Performance comparison of IOH prediction between our HMF and baseline models on two datasets .
 Datasets Sample (s) Model MSE MAE AUC Accuracy (%) Recall (%)
@@
 HMF 165.7575 9.3845 0.6468 69.27 45.87
-Table 3: Ablation study on instance normalization and sequence decomposition in HMF on the CH-OPBP dataset.
-Dataset Model MSE MAE AUC Accuracy (%) Recall (%)
-CH-OPBP HMF (full model) 86.4927 7.2828 0.7413 75.53 70.13
-w/o instance normalization 106.9671 8.5005 0.5891 77.72 20.73
-w/o sequence decomposition 105.7231 8.4964 0.5750 78.10 17.50
+Table 2: Performance comparison of IOH prediction between our HMF and baseline models on two datasets .
 Compared Baselines.
 To evaluate HMF, we compare it with traditional and deep learning-based forecasting models.
-For traditional methods, we use ARIMA Ariyo et al.
-(2014) and Logistic Regression.
+For traditional methods, we use ARIMA Hyndman and Khandakar (2008) and Logistic Regression.
 ARIMA, a statistical time series model, is trained on 0.5% of the test data for efficiency.
@@
 (2018), a hypotensive event is identified when predicted MAP remains below $\theta_{\text{MAP}}=65$ mmHg for at least 1 minute.
-To mitigate the impact of transient fluctuations, a 2-minute skip window Wijnberge et al.
-(2020) is applied, ensuring evaluation focuses on sustained hypotensive trends rather than isolated drops.
+To enable effective early warning and facilitate clinical decision-making, a 2-minute skip window Wijnberge et al.
+(2020) is applied.
 Performance is assessed using accuracy, recall, and Area Under the Curve (AUC), with AUC serving as the primary metric for evaluating the model’s ability to differentiate hypotensive trends.
@@
 Further methodological details, including ground truth labeling and evaluation protocols, are provided in supplement material part.
+Dataset Model MSE MAE AUC Accuracy (%) Recall (%)
+CH-OPBP HMF (Our model) 86.4927 7.2828 0.7413 75.53 70.13
+w/o instance normalization 106.9671 8.5005 0.5891 77.72 20.73
+w/o sequence decomposition 105.7231 8.4964 0.5750 78.10 17.50
+Table 3: Ablation study on instance normalization and sequence decomposition in HMF on the CH-OPBP dataset.
 5.2 Experimental Results
@@
 Compared to LSTM and Transformer, which perform well in some scenarios but face challenges with long sequence dependencies, HMF effectively models temporal dependencies and coupling effects.
-Informer and DLinear, despite their advantages in general forecasting tasks, show limitations in handling intraoperative data, particularly in recall performance.
+Informer and DLinear, despite their advantages in general forecasting tasks, show limitations in handling intraoperative data.
 Overall, HMF maintains stable performance across diverse conditions, demonstrating its potential to enhance IOH prediction in complex clinical settings.
@@
 Integrating instance normalization and sequence decomposition enhances generalization, leading to more effective IOH predictions.
-Table 4: Comparison of transfer and non-transfer learning on individual patients.
 Methods Patient ID MSE MAE AUC
@@
 2 47.6620 6.8317 0.5600
-Table 5: Comparison of transfer and non-transfer learning on elderly and young subgroups.
+Table 4: Comparison of transfer and non-transfer learning on individual patients.
 Methods Feature MSE MAE AUC
@@
 Young 44.1721 5.2447 0.8540
-Performance Analysis w.r.t Transfer Settings.
+Table 5: Comparison of transfer and non-transfer learning on elderly and young subgroups.
+Performance Analysis under Transfer Learning Settings.
 Assessing model transferability across patients and demographic groups is essential for IOH prediction.
@@
 Table 4 presents cross-patient transfer results, where a model trained on one patient is applied to another.
-AUC scores suggest that transfer learning preserves or even enhances predictive performance, indicating the model’s ability to capture shared physiological patterns.
-This reduces reliance on patient-specific data while maintaining accuracy, though individual training may still offer advantages in highly distinct cases.
-Table 5 presents cross-group transfer results between age cohorts.
-The non-transfer model performs better in elderly patients, suggesting that age-specific models better capture physiological complexities.
+AUC scores suggest that transfer learning preserves or enhances predictive performance, indicating the model’s ability to capture shared physiological patterns.
+This reduces reliance on patient-specific data while maintaining accuracy, though individual training may still offer advantages in distinct cases.
+Table 5 presents cross-group results between age cohorts.
+The non-transfer model performs better in elderly patients, suggesting that model better capture physiological complexities.
 In contrast, younger patients benefit from transfer learning, likely due to more homogeneous physiological responses.
-These findings underscore the importance of demographic-specific considerations while demonstrating transfer learning’s potential to improve predictive performance across diverse clinical settings.
+These findings underscore the importance of demographic considerations while demonstrating transfer learning’s potential to improve performance across diverse clinical settings.
 Figure 5: Analysis of MSE and AUC performance for IOH segments across varying $d_{\text{model}}\in\{64,128,256,512,1024\}$.
@@
 Parameter Sensitivity Analysis.
-To optimize key model parameters and enhance the Patch Encoder’s ability to capture temporal dependencies, we conduct a parameter sensitivity analysis on the CH-OPBP dataset, focusing on the model dimension $d_{\text{model}}$ and patch length $S$.
-We first investigate the effect of varying $d_{\text{model}}$ while keeping the context window fixed at 300 and the predicted scope length at 100.
-As shown in Figure 5, different values of $d_{\text{model}}$ significantly influence model performance, with an optimal setting maximizing AUC.
-Subsequently, we examine the impact of varying $S$ while holding $d_{\text{model}}$ constant.
-This analysis is critical, as $S$ directly affects how effectively the model captures local temporal structures within physiological signals.
-The corresponding results, presented in Figure 6, indicate that selecting an appropriate $S$ value plays a crucial role in model performance.
-Based on these analyses, we determine the optimal parameter values that yield the highest AUC scores, ensuring robust model performance.
-The sensitivity of the model’s performance to $d_{\text{model}}$ and $S$ highlights the importance of appropriately configuring the Patch Encoder to accurately model blood pressure dynamics.
-Visualization Case Study Analysis.
-Figure 7 compares MAP predictions across models, demonstrating the effectiveness of the HMF framework.
-Our analysis of the CH-OPBP dataset, using a 15-minute context window and a 3-second sampling rate, consistently shows that HMF outperforms other models in predicting IOH events across various prediction horizons.
-Specifically, for a prediction length of 100, HMF (pink) closely follows actual MAP trends, capturing both rapid fluctuations and long-term variations.
-In contrast, DLinear (purple) and Informer (teal) tend to smooth out critical transitions, while Transformer (yellow) shows occasional delays.
-LSTM (orange) struggles with long-range dependencies, leading to deviations in extended forecasts.
-These results highlight HMF’s advantage in modeling complex temporal patterns for accurate IOH prediction.
-Figure 7: Comparison of MAP prediction across different models on representative test cases.
+To optimize key parameters and enhance the Patch Encoder’s ability to capture temporal dependencies, we perform sensitivity analysis on the CH-OPBP dataset, focusing on model dimension $d_{\text{model}}$ and patch length $S$.
+We first vary $d_{\text{model}}$ with a fixed context window of 300 and prediction length of 100.
+As shown in Figure 5, performance is sensitive to $d_{\text{model}}$, with a clear optimum maximizing AUC.
+We then examine the effect of varying $S$ while keeping $d_{\text{model}}$ constant.
+Since $S$ governs the model’s ability to capture local temporal patterns, its selection is important.
+Results in Figure 6 confirm that appropriate $S$ values significantly affect forecasting accuracy.
+These findings guide the selection of optimal hyperparameters to ensure performance and highlight the importance of configuring the Patch Encoder to effectively model blood pressure dynamics.
 6 Conclusion
-We proposed a novel approach to IOH prediction by reformulating it as a time series forecasting task.
-To address key challenges, we introduced symmetric normalization to mitigate non-stationarity in MAP and SBP series, sequence decomposition to capture complex temporal patterns, and a patch-based Transformer model to enhance representation learning.
-Extensive experiments on real-world datasets validated the effectiveness of our HMF framework, demonstrating its ability to improve early IOH detection.
-This study highlights the potential of time series forecasting in clinical decision support and provides a foundation for future research in intraoperative risk prediction.
-Appendix
-Appendix A Related Work
-A.1 Intraoperative Hypotension Forecasting
-Existing methods for intraoperative hypotension (IOH) prediction predominantly adopt a static modeling approach, extracting handcrafted features from physiological signals rather than directly modeling their temporal dynamics.
-Early studies focused on high-fidelity arterial pressure waveforms, leading to the Hypotension Prediction Index (HPI) Hatib et al.
-[2018b].
-Machine learning approaches, including ensemble methods Cherifa et al.
-[2020] and gradient boosting Kendale et al.
-[2018], integrated preoperative and intraoperative factors but largely treated data as independent samples, overlooking temporal dependencies.
-Deep learning models, such as RNN-based predictors Jeong et al.
-[2019] and attention-based frameworks Lu et al.
-[2023], improved performance by capturing sequential patterns but primarily relied on fixed-length feature representations, failing to fully model non-stationary hemodynamic changes.
-Recent interpretable models Ritter et al.
-[2023] enhanced clinical applicability but remained feature-driven.
-Our work addresses these limitations by formulating IOH prediction as a multivariate time-series forecasting problem, explicitly capturing temporal dependencies and non-stationary dynamics in arterial blood pressure signals.
-A.2 Time Series and Sequence Forecasting
-Time series forecasting is fundamental in domains such as finance, healthcare, and energy Cheng et al.
-[2024b].
-Traditional models like ARIMA Ariyo et al.
-[2014] and exponential smoothing struggle with high-dimensional physiological signals He et al.
-[2023], while deep learning methods, including LSTMs Graves and Graves [2012] and GRUs Dey and Salem [2017], improve long-term dependency modeling.
-Recent Transformer-based models enhance long-sequence forecasting efficiency.
-Informer Zhou et al.
-[2021] optimizes self-attention mechanisms, Autoformer Chen et al.
-[2021] incorporates trend-seasonality decomposition, and FEDformer Zhou et al.
-[2022] leverages frequency domain analysis.
-Lightweight alternatives, such as DLinear Zeng et al.
-[2023] and TSMixer Ekambaram et al.
-[2023], demonstrate competitive performance by preserving temporal structures with simpler architectures.
-GPHT Liu et al.
-[2024] further introduces generative pretraining to improve forecasting transferability across datasets and horizons.
-Despite these advances Cheng et al.
-[2024b]; Zhang et al.
-[2024]; Cheng et al.
-[2024a]; Wang et al.
-[2024b, a], challenges remain in computational efficiency and real-time adaptability for physiological signal forecasting.
-Our work addresses these gaps by integrating multi-factor feature extraction with an efficient Transformer-based approach, improving IOH prediction accuracy in dynamic clinical settings.
-Appendix B Evaluation Details
-B.1 Ground Truth Labeling for IOH Events
-In this study, ground truth labels for intraoperative hypotension (IOH) events were generated based on recorded mean arterial pressure (MAP) values during surgery.
-The labeling process identifies clinically significant hypotensive episodes while minimizing transient fluctuations that may not reflect sustained hemodynamic instability.
-An IOH event is defined as a segment where the MAP value remains below a predefined threshold $\theta_{\text{MAP}}$ for at least $t$ minutes.
-Following clinical guidelines Wesselink et al.
-[2018], we set $\theta_{\text{MAP}}=65$ mmHg, a widely accepted hypotension threshold, and $t=1$ minute to ensure clinical relevance.
-To generate labels, the algorithm scans the MAP time series for each patient, evaluating non-overlapping segments of length $t$.
-For each segment, the maximum MAP value is determined.
-If this value is below $\theta_{\text{MAP}}$, the entire segment is labeled as an IOH event.
-This approach ensures that only sustained hypotensive episodes are captured, reducing false positives from transient MAP drops.
-The labeling process is outlined in Algorithm 2.
-Algorithm 2 Ground Truth Labeling for IOH Events
-Input: $MAP_{\text{actual}}$: recorded MAP series, $t$: minimum duration for an IOH event, $T_{\text{seq}}$: total length of the MAP series.
-Output: $L_{\text{actual}}$: binary label sequence for IOH events.
-1:  Initialize $L_{\text{actual}}[:]=0$
-2:  for $i$ in range(0, $T_{\text{seq}}-t+1$do
-3:     $MAP_{\text{max}}=\max(MAP_{\text{actual}}[i:i+t])$
-4:     if $MAP_{\text{max}}\leq\theta_{\text{MAP}}$ then
-5:        $L_{\text{actual}}[i:i+t]=1$
-6:     end if
-7:  end for
-8:  return $L_{\text{actual}}$
-B.2 Evaluation of Hypotensive Events
-To evaluate IOH detection, we compare the model’s predicted MAP trajectory against actual MAP values, assessing both event identification and duration.
-A predicted IOH event is correct if the forecasted MAP remains below $\theta_{\text{MAP}}$ for at least $t$ minutes.
-Performance is quantified using the following metrics, including precision, recall, and F1-score:
-$\text{Precision}=\frac{\text{TP}}{\text{TP}+\text{FP}}$ (13)
-$\text{Recall}=\frac{\text{TP}}{\text{TP}+\text{FN}}$ (14)
-[…]
-where TP (True Positives) are correctly identified IOH events, FP (False Positives) are incorrect predictions, and FN (False Negatives) are missed hypotensive episodes.
-Additionally, we compute the time-to-detection (TTD), which measures how early the model predicts an IOH event before its actual onset.
-A lower TTD indicates better early-warning capability, which is essential for timely clinical intervention.
-This evaluation framework ensures that our method provides meaningful predictions, distinguishing sustained hypotensive episodes from transient MAP fluctuations while maintaining high predictive accuracy.
-Appendix C Details of Baseline Methods
-To comprehensively evaluate the performance of HMF in early IOH prediction, we benchmarked it against a set of traditional and deep learning-based methods.
-These baselines were selected to cover a range of modeling approaches, from statistical methods to advanced sequence learning techniques.
-C.1 Traditional Methods
-We included two widely used traditional time series forecasting and classification models:
-  • ARIMA Ariyo et al.
-[2014]: The Autoregressive Integrated Moving Average (ARIMA) model was used as a classical statistical forecasting baseline.
-Due to its computational intensity, we trained ARIMA on only 0.5% of the test set, balancing efficiency and representativeness in the comparison.
-  • Logistic Regression: A standard binary classification model, logistic regression was applied to predict IOH events.
-To enhance predictive performance, time series features were extracted using the tsfresh library [^5] https://github.com/blue-yonder/tsfresh, which generated 1,566 handcrafted features, capturing statistical and frequency-domain properties.
-C.2 Deep Learning Methods
-We evaluated several deep learning-based baselines, each known for its effectiveness in modeling sequential dependencies in time series data:
-  • LSTM Graves and Graves [2012]: The Long Short-Term Memory (LSTM) network, a recurrent neural network (RNN) variant, was included due to its capability to capture long-range temporal dependencies.
-Its gating mechanism mitigates vanishing gradient issues, making it suitable for sequential modeling tasks like IOH prediction.
-  • GRU Dey and Salem [2017]: The Gated Recurrent Unit (GRU) is an RNN variant similar to LSTM but with a simplified architecture.
-GRU was included for comparison due to its efficiency and comparable performance in capturing time series patterns while using fewer parameters.
-  • Transformer Vaswani et al.
-[2017]: The Transformer model, with its self-attention mechanism, was included to evaluate the effectiveness of attention-based architectures in IOH prediction.
-By allowing the model to weigh different time steps adaptively, it can capture complex dependencies beyond what RNN-based models achieve.
-  • Informer Zhou et al.
-[2021]: The Informer model extends the Transformer architecture by introducing a ProbSparse self-attention mechanism, reducing computational cost while preserving long-range dependency modeling.
-Its efficiency makes it well-suited for long-sequence forecasting tasks.
-  • DLinear Zeng et al.
-[2023]: A lightweight linear model for time series forecasting, DLinear models trend and seasonal components separately.
-Despite its simplicity, it has demonstrated competitive performance against complex deep learning models, providing a strong baseline.
-  • TSMixer Ekambaram et al.
-[2023]: A recent MLP-based model, TSMixer improves time series forecasting by using a structured mixing approach for temporal dependencies.
-We included it to assess the performance of purely feedforward architectures compared to sequential and attention-based models.
-These baselines provide a diverse comparison set, ranging from traditional statistical methods to state-of-the-art deep learning architectures.
-By including both simple and complex models, we ensure a rigorous evaluation of HMF’s effectiveness in IOH prediction.
-Table 6: Performance Comparison of Different Methods Using The CH-OPBP Dataset.
-Model Sample(s) Pred MSE MAE AUC Accuracy (%) Recall (%)
-Arima 1 300 97.1580 7.5501 0.6071 90.34 23.33
-600 118.0238 8.8539 0.6022 75.15 25.53
-900 175.6289 10.1539 0.5795 66.48 23.13
-3 100 26.9210 4.0537 0.6625 86.60 38.46
-200 155.1008 10.0347 0.5870 77.27 22.22
-300 156.7626 10.3933 0.5288 63.03 16.28
-LR 1 300 0.5000 86.06 0.00
-600 0.6777 62.13 81.90
-900 0.3386 35.96 23.39
-3 100 0.5977 79.84 32.50
-200 0.7002 72.00 65.18
-300 0.7342 75.01 65.79
-LSTM 1 300 98.2363 7.9263 0.5660 90.65 14.57
-600 122.3699 9.4033 0.5000 73.49 0.00
-900 133.7677 9.9659 0.5226 64.47 5.18
-3 100 104.2937 8.9187 0.5000 90.51 0.00
-200 127.8037 10.0756 0.5000 73.48 0.00
-300 141.1666 10.6498 0.5000 62.94 0.00
-Transformer 1 300 106.0975 8.2478 0.6194 89.99 27.32
-600 126.0585 9.6861 0.5804 75.94 19.94
-900 148.2355 10.2089 0.5758 67.32 20.04
-3 100 93.5870 7.4800 0.6328 89.51 30.91
-200 112.6674 9.0762 0.5406 75.06 9.34
-300 108.3092 8.4761 0.6177 70.22 29.11
-Informer 1 300 102.4836 8.2318 0.5663 90.41 14.95
-600 92.3734 7.6706 0.6952 79.20 48.91
-900 116.2513 8.4248 0.6741 73.93 42.27
-3 100 104.9414 8.1057 0.6065 90.11 24.28
-200 108.3698 8.7226 0.5934 77.23 21.25
-300 119.7905 8.3365 0.6835 74.24 45.59
-DLinear 1 300 107.2025 8.2388 0.5468 90.39 10.61
-600 127.4571 9.6158 0.5316 74.75 7.20
-900 142.3762 10.1150 0.5208 64.33 4.82
-3 100 106.1303 8.2176 0.5587 90.27 13.42
-200 123.7473 9.4676 0.5711 76.31 16.21
-300 141.7650 10.2002 0.5271 64.72 6.30
-HMF 1 300 73.7196 6.8555 0.7702 76.48 77.70
-600 89.8372 7.6161 0.7230 75.30 65.91
-900 116.2464 8.2751 0.7125 74.08 60.33
-3 100 77.0390 6.8708 0.7801 77.72 78.38
-200 88.6844 7.4293 0.7245 75.32 66.34
-300 93.7548 7.5482 0.7194 73.55 65.68
-Table 7: Performance Comparison of Different Methods Using The VitalDB Dataset.
-Model Sample(s) Pred MSE MAE AUC Accuracy(%) Recall(%)
-Arima 3 100 193.3889 10.7640 0.5397 83.10 10.42
-200 268.3629 13.7636 0.5073 60.27 5.48
-300 310.3584 14.8105 0.5280 53.20 9.68
-LR 3 100 0.5094 78.52 2.11
-200 0.5877 61.80 37.50
-300 0.5814 58.41 60.80
-LSTM 3 100 157.8527 10.6642 0.5000 84.35 0.00
-200 204.0663 12.3266 0.5000 61.35 0.00
-300 204.2178 12.5930 0.5000 48.58 0.00
-Transformer 3 100 162.8149 10.6928 0.5022 84.36 0.54
-200 162.6087 10.8778 0.5006 61.40 0.13
-300 150.6858 10.4997 0.5091 49.50 2.12
-Informer 3 100 164.9233 11.0281 0.5000 84.35 0.00
-200 152.0211 10.7310 0.5000 61.35 0.00
-300 159.4176 10.9371 0.5008 48.66 0.16
-DLinear 3 100 173.4594 11.1419 0.5097 84.33 2.41
-200 175.6332 11.4898 0.5077 61.85 1.97
-300 176.2505 11.8588 0.5049 49.08 1.19
-HMF 3 100 149.5450 8.7662 0.6819 78.29 53.50
-200 165.2017 9.3708 0.6370 68.01 44.74
-300 182.5259 10.0166 0.6215 61.50 39.38
-Appendix D Appendix D: Full Results Report
-Tables 6 and 7 present the full experimental results comparing the performance of different baseline methods and our proposed HMF model on the CH-OPBP and VitalDB datasets.
-These tables provide a detailed evaluation across multiple metrics, including mean squared error (MSE), mean absolute error (MAE), area under the curve (AUC), accuracy, and recall, under different prediction horizons (e.g., 100, 200, and 300 sample steps).
-The results highlight the effectiveness of HMF in capturing intraoperative hypotension (IOH) dynamics compared to traditional and deep learning-based approaches.
-Notably, HMF demonstrates superior performance in key metrics such as AUC and recall, indicating its ability to better identify hypotensive events while maintaining predictive accuracy.
+In this paper, we proposed a novel approach to intraoperative hypotension (IOH) prediction by reformulating it as a dynamic sequence forecasting task, enabling proactive and continuous risk assessment.
+To address key challenges inherent in modeling arterial blood pressure dynamics, we introduced symmetric normalization to mitigate non-stationarity in mean arterial pressure (MAP) and systolic blood pressure (SBP) series, sequence decomposition to better capture complex temporal patterns, and a patch-based Transformer model to enhance representation learning and forecasting accuracy.
+Extensive experiments on real-world intraoperative datasets validated the effectiveness and robustness of our HMF framework, demonstrating its strong performance in improving early IOH forecasting under various clinical conditions.
+Acknowledgements
+This research was supported by grants from the grants of Provincial Natural Science Foundation of Anhui Province (No.2408085QF193), USTC Research Funds of the Double First-Class Initiative (No.
+YD2150002501), the Fundamental Research Funds for the Central Universities of China(Grant No.
+PA2024GDSK0112, No.
+WK2150110032), Anhui Provincial Natural Science Foundation (No.
+2308085MG226), the Key Technologies R&D Program of Anhui Province (No.
+202423k09020039).